A positive charge of 4.00 μC is fixed in place. From a distance of 3.80 cm a particle of mass 5.50 g and charge +3.30 μC is fired with an initial speed of 66.0 m/s directly toward the fixed charge. How close to the fixed charge does the particle get before it comes to rest and starts traveling away?
Answer #1

using energy conservation

KE initally = change in PE of system

0.5 m v^2 = k q1 q2 * ( 1/ r1 - 1/ r2)

0.5* 5.5*10^-3* 66^2 = 9*10^9* 3.3* 4*10^-12* ( 1/ r1 - 1/ 0.038)

r1 = 0.00786 m

=========

Comment in case any doubt, will reply for sure. Goodluck